Question 244 pts

Professor Dean Winchester wanted to investigate whether people who believe in vampires are likely to buy garlic. People registered on creaturesofthenight.com were asked how many bulbs of garlic they bought last month. He obtained the following data from 4 participants: 10, 8, 10, 12.

Which is the BEST measure of central tendency for this data:

Mean
Median
Mode
zscore
Answer #1

from 4 participants: 8,10, 10, 12 (in ascending order)

From these data:

Mean(xbar)=(10+8+10+12)/4=10

Median=(10+10)/2=10

Mode=10 (it has highest frequency)

Since mode=mean=median

Hence we Can assume it normally distributed so

Z-score will be the BEST measure of central tendency for this data